# DHP: Discrete Hierarchical Planning for HRL Agents

### v2

@@ Abstract
Hierarchical Reinforcement Learning (HRL) agents often struggle with long-horizon visual planning due to their reliance on error-prone distance metrics. We propose Discrete Hierarchical Planning (DHP), a method that replaces continuous distance estimates with discrete reachability checks to evaluate subgoal feasibility. DHP recursively constructs tree-structured plans by decomposing long-term goals into sequences of simpler subtasks, using a novel advantage estimation strategy that inherently rewards shorter plans and generalizes beyond training depths. In addition, to address the data efficiency challenge, we introduce an exploration strategy that generates targeted training examples for the planning modules without needing expert data. Experiments in 25-room navigation environments demonstrate $100\%$ success rate (vs $82\%$ baseline) and $73$-step average episode length (vs $158$-step baseline). The method also generalizes to momentum-based control tasks and requires only $\log N$ steps for replanning. Theoretical analysis and ablations validate our design choices.
@@ 1 Introduction
Hierarchical planning enables agents to solve complex tasks through recursive decomposition [8], but existing approaches face fundamental limitations in long-horizon visual domains. While methods using temporal distance metrics [27, 2] or graph search [10] have shown promise, their reliance on precise continuous distance estimation creates two key challenges:
  • Coupled Learning Dynamics: Distance estimates depend on the current policy’s quality, where suboptimal policies produce misleading distances and practical implementations may require arbitrary distance cutoffs [2, 10].
@@
We address these issues by reformulating hierarchical planning through discrete reachability – a paradigm shift from "How far?" to "Can I get there?". Our method (DHP) evaluates plan feasibility through binary reachability checks rather than continuous distance minimization. This approach builds on two insights: local state transitions are easier to model than global distance metrics, and reachability naturally handles disconnected states through $0/1$ signaling.
@@
  • A reachability-based discrete reward scheme for planning (Sec. 2.3.2, Fig. 8(d))
  • An advantage estimation algorithm for tree-structured trajectories (Sec. 2.3.3, Fig. 8(b)).
@@
Empirical results confirm these design choices. On the 25-room navigation benchmark, DHP achieves perfect success rates ($100\%$ vs $82\%$) with 2.1x shorter trajectories than state-of-the-art (Sec. 3.1). Our ablation studies demonstrate that both the reachability paradigm and the novel advantage estimation contribute significantly to these gains (Section 3.2). We also test our method in a momentum-based RoboYoga environment for generalizability. The code is available at: <URL redacted>.
@@ 2 Discrete Hierarchical Planning (DHP)
In the context of Markovian Decision Processes (MDP), a Reinforcement Learning task can be imagined as an agent transitioning through states $s_{t}$ using actions $a_{t}$. Our task is to find the shortest path between any two given states $(s_{t},s_{g})$. To do this, we first use an explorer to collect a dataset of useful trajectories possible within the environment. Then we learn a planning policy that learns to predict subgoals that decompose the initial task into two simpler sub-tasks. The recursive application of the policy breaks the subtasks further till subtasks directly manageable by the worker are found.
@@ 2.1 Agent Architecture
We use the base architecture from the Director [14], which consists of three modules: world-model, worker, and manager. World model is implemented using the Recurrent State Space Module (RSSM) [13] that learns state representations $s_{t}$ using a sequence of image observations $o_{t}$. The worker policy $\pi_{W}$ learns to output environmental actions $a_{t}$ to reach nearby states $s_{wg}$. The manager $\pi_{M}$ is a higher-level policy that learns to output desirable sub-goal states $s_{wg}$ for the worker (updated every $K$ steps) in the context of an external task and the exploratory objective jointly. We use our explorer policy $\pi_{E}$ and the planning policy $\pi_{P}$ as the manager. Figure 1(a) shows the agent architecture during inference using $\pi_{P}$. The worker and the world-model are trained using the default objectives from the Director (for more details see Appendix B). First, we describe our planning policy and then derive an exploratory strategy fit for it.
@@ 2.2 Planning Policy
The Planning policy $\pi_{P}$ is a goal-conditioned policy that takes the current and goal state as inputs to yield a subgoal. However, predicting directly in the continuous state space leads to the problem of high-dimensional continuous control [14, 27]. To reduce the search space for the planning policy, we train a discrete Conditional Variational AutoEncoder (CVAE) that learns to predict midway states $s_{t+q/2}$ given an initial and a final state $(s_{t},s_{t+q})$. We call this module Goal-Conditioned State Recall (GCSR). It consists of an Encoder and a Decoder (Fig. 1(b)). The Encoder takes the initial, midway, and final states $(s_{t},s_{t+q/2},s_{t+q})$ as input to output a distribution over a latent variable $\text{Enc}_{G}(z|s_{t},s_{t+q/2},s_{t+q})$. The Decoder uses the states $(s_{t},s_{t+q})$ and a sample from the latent distribution $z\sim\text{Enc}_{G}(z)$ to predict the midway state $\text{Dec}_{G}(s_{t},s_{t+q},z)\rightarrow\hat{s}_{t+q/2}$. The module is optimized using the explorer collected data to minimize the ELBO objective (Eq. 1). A mixture of categoricals $(4\times 4)$ as the prior latent distribution $p_{G}(z)$ for all our experiments. Triplets at multiple temporal resolutions $q\in Q$ are extracted, allowing the agent to function at all temporal resolutions. For minimum overlap between resolutions, we use exponentially increasing temporal resolutions, $Q=\{2K,4K,8K,...\}$.
@@
Figure 2: The figure illustrates the plan unrolling process during the training and inference phases. (a) During Training, the initial subtask $(s_{t},s_{g})$ is recursively decomposed into smaller tasks by midway subgoal prediction to generate a subtask tree. The lowest level nodes represent the simplest decomposition of the initial task as: $(s_{t},s_{g})\rightarrow(s_{t},s_{3},s_{1},s_{4},s_{0},s_{5},s_{2},s_{6},s_{g})$. (b) During Inference, only the first branch of the tree is unrolled. Here, the agent is tasked to get to $s_{g}$. So it first divides the task $(s_{t},s_{g})$ into two chunks by inserting $s_{0}$. Then it proceeds to divide the subtask $(s_{t},s_{0})$ by inserting $s_{1}$ and ignoring the second part $(s_{0},s_{g})$. The recursive division continues till the first subgoal reachable in $K$ steps is found. This results in a stack of subgoals shown at the bottom.
@@ 2.3 Planning Policy Optimization
The planning policy is optimized as a Soft-Actor-Critic (SAC) [12] in three steps: construct plans between random state pairs (Sec. 2.3.1), plan evaluation using discrete rewards and our novel advantage estimation method (Sec. 2.3.2,2.3.3), and policy updates using policy gradients (Sec. 2.3.4).
@@ 2.3.1 Plan Unrolling
Given the initial and final states $(s_{t},s_{g})$, a subgoal generation method predicts an intermediate subgoal $s_{0}$ that breaks the task into two simpler subtasks $(s_{t},s_{0})$ and $(s_{0},s_{g})$. The recursive application of the subgoal operator breaks the initial task, leading to a tree of subtasks $\tau$. The root node $n_{0}$ represents the original task $(s_{t},s_{g})$, and each remaining node $n_{i}$ in the tree $\tau$ represents a sub-task. At each node $n_{i}$, the policy predicts a subgoal as: $s_{i}=\text{Dec}_{G}(n_{i,0},n_{i,1},z)$ where $z\sim\pi_{P}(z_{i}|n_{i})$. The preorder traversal of the subtask tree of depth $D$ can be written as $n_{0},n_{1},n_{2},...,n_{2^{D+1}-2}$. Figure 2(a) shows an example unrolled tree. The lowest-level nodes show the smallest decompositions of the task under the current planning policy $\pi_{P}$.
Inference: Unlike traditional methods for hierarchical planning (eg, cross-entropic methods (CEM) [29, 22]), which require unrolling multiple full trees followed by evaluation at runtime, our method does not require full tree expansion. A learned policy always predicts the best estimated subgoal by default. This also means that we do not need to unroll the second subtask further at each node because only the first reachable subgoal is required (Fig. 2(b)).
@@ 2.3.2 Discrete Rewards Scheme
We want to encourage trees that end in subtasks manageable by the worker. A subtask is worker-manageable if the node goal $n_{i,1}$ is reachable by the worker from the node initial state $n_{i,0}$. Since learning modules inside other learned modules can compound errors, we use a more straightforward method to check reachability. We simulate the worker for $K$ steps using RSSM imagination, initialized at $n_{i,0}$ with goal $n_{i,1}$. If the cosine_max similarity measure between the worker’s final state $s_{t,i}$ and the assigned goal state is above a threshold $\Delta_{R}$, the node is marked as terminal (Eq. 2). The terminal nodes do not need further expansion, which is different from the word’s usual meaning in the context of trees. And the lowest layer nodes in a finite depth unrolled tree are called truncated nodes. Computing the terminal array $T_{i}$ allows supporting imperfect trees with branches terminating at different depths. A plan is considered valid if all its branches end in terminal nodes. The policy is rewarded $1$ at terminal nodes and 00 otherwise (Eq. 3). A discrete reward scheme allows for optimization that increases the likelihood of valid plans over distance-based approaches that gradually optimize the policy to reduce path length.
@@ 2.3.3 Return Estimation for Trees
Taking inspiration from the standard discounted return estimation for linear trajectories (sequence of states) [31], we propose an approach for tree trajectories. Returns for a linear trajectory are computed as the reward received at the next step and discounted rewards thereafter, $G_{t}=R_{t+1}+\gamma G_{t+1}$. Similarly, the return estimation for trees is the minimum discounted return from the child nodes. Given a tree trajectory $\tau$, the Monte-Carlo return, the $1$-step return, and the lambda return for each non-terminal node as can be written as:
@@
$G_{i}^{0}=(1-T_{i})\cdot\min(R_{2i+1}+\gamma v_{P}(n_{2i+1}),R_{2i+2}+\gamma v_{P}(s_{2i+2}))$ (5)
$G_{i}^{\lambda}=(1-T_{i})\cdot\min(R_{2i+1}+\gamma((1-\lambda)v_{P}(n_{2i+1})+\lambda G_{2i+1}^{\lambda}),R_{2i+2}+\gamma((1-\lambda)v_{P}(s_{2i+2})+\lambda G_{2i+2}^{\lambda}))$ (6)
All branches should end in terminal nodes to score a high return with the above formulation. Also, since the discount factor diminishes the return with each additional depth, the agent can score higher when the constructed tree is shallow (less maximum depth). This characteristic is similar to linear trajectories, where the return is higher for shorter paths to the goal [33].
Fig. 3 illustrates example return evaluations for Monte-Carlo returns and $n$-step truncated returns that use value estimates to replace rewards at the non-terminal truncated nodes. The $n$-step returns allow for generalization beyond the maximum unrolled depth $D$ (Sec. A.2.2). Thus, the tree can be unrolled for higher depths $D_{I}$ during inference. We show that the Bellman operators for the above returns are contractions and repeated applications cause the value function $v^{\pi}_{P}$ to approach a stationary $v^{*}$ (Sec. A.4). We explore how the return penalizes maximum tree depth and encourages balanced trees in Sec. A.2.3, implying that the policy breaks tasks roughly halfway.
@@
Figure 3: Example return estimations for an imperfect tree (node indices at top-left). The dash-bordered cells indicate terminal nodes where the policy receives a $1$ reward and the branch terminates. While one of the branches terminates early $(i=4)$, one does not for the unrolled depth $(i=11)$. (a) Since we compute the return as the $\min$ of the child nodes, the Monte-Carlo return at the root node is 00 in this case. However, there is still a positive learning signal induced at nodes $(i=1,3,6)$. (b) Using the critic as a baseline for computing $n$-step returns. The $n$-step returns allow bootstrapping by substituting the reward with value estimates $v$ at the truncated node $(i=11)$. This induces a learning signal at the root node even if the plan is incomplete for the unrolled depth.
@@ 2.3.4 Policy Gradients
Using tree return estimates (we use $n$-step lambda returns in all cases), we derive the policy gradients for the planning policy as (Eq. 7, proof in Sec. A.1). We show that if the function $G^{\lambda}_{i}$ is replaced by a function independent of the policy actions, the expectation reduces to 00, implying that we can use the value function as a baseline for variance reduction (Th. A.2). With policy gradients and an entropy term, to encourage random exploration before convergence, we construct the loss function for the actor $\pi_{P}$ and critic $v_{P}$ as (sum over all nodes except the terminal and truncated nodes):

### v3

@@ Abstract
Hierarchical Reinforcement Learning (HRL) agents often struggle with long-horizon visual planning due to their reliance on error-prone distance metrics. We propose Discrete Hierarchical Planning (DHP), a method that replaces continuous distance estimates with discrete reachability checks to evaluate subgoal feasibility. DHP recursively constructs tree-structured plans by decomposing long-term goals into sequences of simpler subtasks, using a novel advantage estimation strategy that inherently rewards shorter plans and generalizes beyond training depths. In addition, to address the data efficiency challenge, we introduce an exploration strategy that generates targeted training examples for the planning modules without needing expert data. Experiments in 25-room navigation environments demonstrate a $100\%$ success rate (vs. $90\%$ baseline). We also present an offline variant that achieves state-of-the-art results on OGBench benchmarks, with up to $71\%$ absolute gains on giant HumanoidMaze tasks, demonstrating our core contributions are architecture-agnostic. The method also generalizes to momentum-based control tasks and requires only $\log N$ steps for replanning. Theoretical analysis and ablations validate our design choices.
@@ 1 Introduction
Hierarchical planning enables agents to solve complex tasks through recursive decomposition [dietterich2000hierarchical], but existing approaches face fundamental limitations in long-horizon visual domains. While methods using temporal distance metrics [pertsch2020long, ao2021co] or graph search [eysenbach2019search] have shown promise, their reliance on precise continuous distance estimation creates two key challenges:
  • Coupled Learning Dynamics: Distance estimates depend on the current policy’s quality, where suboptimal policies produce misleading distances and practical implementations may require arbitrary distance cutoffs [ao2021co, eysenbach2019search].
@@
We address these issues by reformulating hierarchical planning through discrete reachability – a paradigm shift from "How far?" to "Can I get there?". Our method (DHP) evaluates plan feasibility through binary reachability checks rather than continuous distance minimization. This approach builds on two insights: local state transitions are easier to model than global distance metrics, and reachability naturally handles disconnected states through $0/1$ signaling. To demonstrate the broad applicability of our approach, we develop both online and offline variants. The online version (Sections 2) demonstrates the complete system with world model-based imagination and intrinsic exploration. The offline variant (Section F) isolates our core contributions—discrete reachability and tree-structured returns—showing they transfer successfully to different architectural choices and more complex environments.
@@
  • A reachability-based discrete reward scheme for planning. (Sec. 2.3.2, Fig. 8(d)).
  • An advantage estimation algorithm for tree-structured plans (Sec. 2.3.3, Fig. 8(b)).
@@
  • An offline variant demonstrating architecture-agnostic applicability (Sec. F).
Empirical results confirm these design choices. On the 25-room navigation benchmark, DHP achieves perfect success rates ($100\%$ vs $90\%$) with shorter path lengths (Sec. 3.1). Our offline variant achieves substantial improvements on the more challenging OGBench environments, with up to $71.2\%$ absolute gains on HumanoidMaze-Giant (Sec. 3.3). Our ablation studies demonstrate that both the reachability paradigm and the novel advantage estimation contribute significantly to these gains (Section 3.2). We also test our method in a momentum-based RoboYoga environment for generalizability. Video: https://sites.google.com/view/dhp-video/home.
@@ 2 Discrete Hierarchical Planning (DHP)
In the context of Markovian Decision Processes (MDP), a Reinforcement Learning task can be imagined as an agent transitioning through states $s_{t}$ using actions $a_{t}$. Our task is to find the shortest path between any two given states $(s_{t},s_{g})$. To do this, we first use an explorer to collect a dataset of useful trajectories possible within the environment. Then we learn a planning policy that learns to predict subgoals, decomposing the initial task into two simpler subtasks. The recursive application of the policy breaks the subtasks further till subtasks directly manageable by the worker are found.
@@ 2.1 Agent Architecture
We use the base architecture from the Director [hafner2022deep] as it has been observed to provide a practical method for learning hierarchical behaviors directly from pixels. It consists of three modules: world-model, worker, and manager. The world model is implemented using the Recurrent State Space Module (RSSM) [hafner2019learning], which learns state representations $s_{t}$ using a sequence of image observations $o_{t}$. The worker policy $\pi_{W}$ learns to output environmental actions $a_{t}$ to reach nearby states $s_{wg}$. The manager $\pi_{M}$ is a higher-level policy that learns to output desirable sub-goal states $s_{wg}$ for the worker (updated every $K$ steps) in the context of an external task and the exploratory objective jointly. We replace the manager with our explorer policy $\pi_{E}$ and the planning policy $\pi_{P}$ as required. Figure 1(a) shows the agent architecture during inference using $\pi_{P}$. The worker and the world-model are trained using the default objectives from the Director (for more details see Appendix B). First, we describe our planning policy and then derive an exploratory strategy fit for it.
@@ 2.2 Planning Policy
The Planning policy $\pi_{P}$ is a goal-conditioned policy that takes the current and goal state as inputs to yield a subgoal. However, predicting directly in the continuous state space leads to the problem of high-dimensional continuous control [hafner2022deep, pertsch2020long]. To reduce the search space for the planning policy, we train a discrete Conditional Variational AutoEncoder (CVAE) that learns to predict midway states $s_{t+q/2}$ given an initial and a final state $(s_{t},s_{t+q})$. We refer to this module as Goal-Conditioned State Recall (GCSR). It consists of an Encoder and a Decoder (Fig. 1(b)). The Encoder takes the initial, midway, and final states $(s_{t},s_{t+q/2},s_{t+q})$ as input to output a distribution over a latent variable $\text{Enc}_{G}(z|s_{t},s_{t+q/2},s_{t+q})$. The Decoder uses the states $(s_{t},s_{t+q})$ and a sample from the latent distribution $z\sim\text{Enc}_{G}(z)$ to predict the midway state $\text{Dec}_{G}(s_{t},s_{t+q},z)\rightarrow\hat{s}_{t+q/2}$. The module is optimized using the data collected by the explorer to minimize the ELBO objective (Eq. 1). A mixture of categoricals $(4\times 4)$ is used as the prior latent distribution $p_{G}(z)$ for all our experiments. Triplets at multiple temporal resolutions $q\in Q$ are extracted, allowing the agent to function at all temporal resolutions. For minimizing overlap between resolutions, we use exponentially increasing temporal resolutions, $Q=\{2K,4K,8K,$...\}.
@@
Figure 2: The figure illustrates the plan unrolling process during the training and inference phases. (a) During Training, the initial task $(s_{t},s_{g})$ is recursively decomposed into smaller tasks by midway subgoal prediction to generate a subtask tree. The lowest level nodes represent the simplest decomposition of the initial task as: $(s_{t},s_{g})\rightarrow(s_{t},s_{3},s_{1},s_{4},s_{0},s_{5},s_{2},s_{6},s_{g})$. (b) During Inference, only the first branch of the tree is unrolled. Here, the agent is tasked with reaching $s_{g}$. So it first divides the task $(s_{t},s_{g})$ into two chunks by inserting $s_{0}$. Then it proceeds to divide the subtask $(s_{t},s_{0})$ by inserting $s_{1}$ and ignoring the second part $(s_{0},s_{g})$. The recursive division continues till the first subgoal reachable in $K$ steps is found. This results in a stack of subgoals shown at the bottom.
@@ 2.3 Planning Policy Optimization
The planning policy is optimized as a Soft-Actor-Critic (SAC) [haarnoja2018soft] in three steps: construct plans between random state pairs (Sec. 2.3.1), plan evaluation using discrete rewards and our novel advantage estimation method (Sec. 2.3.2,2.3.3), and policy updates using policy gradients (Sec. 2.3.4).
@@ 2.3.1 Plan Unrolling
Given the initial and final states $(s_{t},s_{g})$, subgoal generation methods predict an intermediate subgoal $s_{0}$ that breaks the task into two simpler subtasks $(s_{t},s_{0})$ and $(s_{0},s_{g})$. The recursive application of the subgoal operator further breaks the task, leading to a tree of subtasks $\tau$. The root node $n_{0}$ represents the original task $(s_{t},s_{g})$, and each remaining node $n_{i}$ in the tree $\tau$ represents a sub-task. At each node $n_{i}$, the policy predicts a subgoal as: $s_{i}=\text{Dec}_{G}(n_{i,0},n_{i,1},z)$ where $z\sim\pi_{P}(z_{i}|n_{i})$. The preorder traversal of the subtask tree of depth $D$ can be written as $n_{0},n_{1},n_{2},...,n_{2^{D+1}-2}$. Figure 2(a) shows an example unrolled tree. The lowest-level nodes show the smallest decompositions of the task under the current planning policy $\pi_{P}$.
Inference: Unlike traditional methods for hierarchical planning (eg, cross-entropic methods (CEM) [rubinstein2004cross, nagabandi2018neural]), which require unrolling multiple full trees followed by evaluation at runtime, our method does not require full tree expansion. A learned policy always predicts the best estimated subgoal by default. Thus, we can unroll only the leftmost branch, as only the first reachable subgoal is required (Fig. 2(b)). Efficient planning allows the agent to re-plan at every goal refresh step ($K$), thereby tackling dynamic and stochastic environments.
@@ 2.3.2 Discrete Rewards Scheme
We want to encourage trees that end in subtasks manageable by the worker. A subtask is worker-manageable if the node goal $n_{i,1}$ is reachable by the worker from the node initial state $n_{i,0}$. Since learning modules inside other learned modules can compound errors, we use a more straightforward method to check reachability. We simulate the worker for $K$ steps using RSSM imagination, initialized at $n_{i,0}$ with goal $n_{i,1}$. If the cosine_max similarity measure between the worker’s final state $s_{t,i}$ and the assigned goal state is above a threshold $\Delta_{R}$, the node is marked as terminal (Eq. 2). The terminal nodes do not need further expansion, which is different from the word’s usual meaning in the context of trees. The lowest layer non-terminal nodes in a finite-depth unrolled tree are called truncated nodes. Computing the terminal array $T_{i}$ allows supporting imperfect trees with branches terminating at different depths. A plan is considered valid if all its branches end in terminal nodes. The policy is rewarded $1$ at terminal nodes and 0 otherwise (Eq. 3). A discrete reward scheme enables optimization that increases the likelihood of valid plans compared to distance-based approaches, which gradually optimize the policy to reduce path length.
@@ 2.3.3 Return Estimation for Trees
Taking inspiration from the standard discounted return estimation for linear trajectories (sequence of states) [sutton2018reinforcement], we propose an approach for tree trajectories. Returns for a linear trajectory are computed as the reward received at the next step and discounted rewards thereafter, $G_{t}=R_{t+1}+\gamma G_{t+1}$. Similarly, the return estimate for trees is the minimum discounted return from the child nodes. Given a tree trajectory $\tau$, the Monte-Carlo return (Eq. 4), the $1$-step return (Eq. 5), and the lambda return (Eq. 6) for each non-terminal node as can be written as:
@@
$G_{i}^{0}=(1-T_{i})\cdot\min(R_{2i+1}+\gamma v_{P}(n_{2i+1}),R_{2i+2}+\gamma v_{P}(n_{2i+2}))$ (5)
$G_{i}^{\lambda}=(1-T_{i})\cdot\min(R_{2i+1}+\gamma((1-\lambda)v_{P}(n_{2i+1})+\lambda G_{2i+1}^{\lambda}),R_{2i+2}+\gamma((1-\lambda)v_{P}(n_{2i+2})+\lambda G_{2i+2}^{\lambda}))$ (6)
Intuition: The min-child formulation ensures a plan is only as valuable as its weakest subtask. If either branch fails to reach a reachable subgoal, the entire plan receives low return. This naturally implements our discrete reachability criterion: good plans must decompose the task into two reachable subtasks. Unlike sum-based formulations (which tolerate one bad child), or average-based formulations (which dilute failure signals), our approach provides clear credit assignment to the bottleneck in the plan.
All branches should end in terminal nodes to score a high return with the above formulation. Additionally, since the discount factor diminishes the return with each additional depth, the agent can score higher when the constructed tree is shallow (less maximum depth). This characteristic is similar to linear trajectories, where the return is higher for shorter paths to the goal [tamar2016value].
Fig. 3 illustrates example return evaluations for Monte-Carlo returns and $n$-step truncated returns that use value estimates to replace rewards at the non-terminal truncated nodes. The $n$-step returns allow for generalization beyond the maximum unrolled depth $D$ (Sec. A.3.2). Thus, the tree can be unrolled for higher depths $D_{I}$ during inference. We show that the Bellman operators for the above returns are contractions and repeated applications cause the value function $v^{\pi}_{P}$ to approach a stationary $v^{*}$ (Sec. A.4). We explore how the return penalizes maximum tree depth and encourages balanced trees (Sec. A.3.3), implying that the optimal policy inherently breaks tasks roughly halfway.
@@
Figure 3: Example return estimations for an imperfect tree (node indices at top-left). The dash-bordered cells indicate terminal nodes where the policy receives a $1$ reward and the branch terminates. While one of the branches terminates early $(i=4)$, one does not for the unrolled depth $(i=11)$. (a) Since we compute the return as the $\min$ of the child nodes, the Monte-Carlo return at the root node is 0 in this case. However, a positive learning signal is still induced at nodes $(i=1,3,6)$. (b) Using the critic as a baseline for computing $n$-step returns. The $n$-step returns allow bootstrapping by substituting the reward with value estimates $v$ at the truncated node $(i=11)$. This induces a learning signal at the root node even if the plan is incomplete for the unrolled depth.
@@ 2.3.4 Policy Gradients
Using tree return estimates (we use $n$-step lambda returns in all cases), we derive the policy gradients for the planning policy as (Eq. 7, proof in Sec. A.2). We show that if the function $G^{\lambda}_{i}$ is replaced by a function independent of the policy actions, the expectation reduces to 0, implying that we can use the value function as a baseline for variance reduction (Th. A.2). With policy gradients and an entropy term, to encourage random exploration before convergence, we construct the loss function for the actor $\pi_{P}$ and critic $v_{P}$ as (sum over all nodes except the terminal and truncated nodes):